### v1

# Bilevel ZOFO: Bridging Parameter-Efficient and Zeroth-Order Techniques for Efficient LLM Fine-Tuning and Meta-Training

Abstract
Fine-tuning pre-trained Large Language Models (LLMs) for downstream tasks using First-Order (FO) optimizers presents significant computational challenges. Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed to address these challenges by freezing most model parameters and training only a small subset. While PEFT is efficient, it may not outperform full fine-tuning when high task-specific performance is required. Zeroth-Order (ZO) methods offer an alternative for fine-tuning the entire pre-trained model by approximating gradients using only the forward pass, thus eliminating the computational burden of back-propagation in first-order methods. However, when implementing ZO methods, a hard prompt is crucial, and relying on simple, fixed hard prompts may not be optimal. In this paper, we propose a bilevel optimization framework that complements ZO methods with PEFT to mitigate sensitivity to hard prompts while efficiently and effectively fine-tuning LLMs. Our Bilevel ZOFO (Zeroth-Order-First-Order) method employs a double-loop optimization strategy, where only the gradient of the PEFT model and the forward pass of the base model are required. We provide convergence guarantees for Bilevel ZOFO. Empirically, we demonstrate that Bilevel ZOFO outperforms both PEFT and ZO methods in single-task settings while maintaining similar memory efficiency. Additionally, we show its strong potential for multitask learning. Compared to current first-order meta-training algorithms for multitask learning, our method has significantly lower computational demands while maintaining or improving performance.
Machine Learning, ICML
1 Introduction
Fine-tuning pretrained Large Language Models (LLMs) has become a standard approach for downstream tasks. Traditional first-order (FO) optimizers like Adam (Kingma & Ba, 2015), commonly used for this process, rely on backpropagation. However, as highlighted in Malladi et al. (2023), computing gradients for LLMs can require up to 12 times the memory needed for inference. This scaling challenge becomes even more pronounced as models grow larger, imposing significant memory demands and complicating the fine-tuning process, especially in resource-constrained environments.
To address these computational challenges, Parameter-Efficient Fine-Tuning (PEFT) methods have been developed. These techniques freeze most of the model’s parameters and train only a small subset, significantly reducing both memory and computational overhead. Popular PEFT approaches include prompt tuning, LoRA fine-tuning, and prefix tuning. Prompt tuning (Lester et al., 2021; Qin & Eisner, 2021; Yu et al., 2023) optimizes continuous prompt vectors that are concatenated with the input embeddings, while prefix tuning (Li & Liang, 2021) introduces learnable prefix tokens that serve as conditioning variables at each transformer layer. LoRA (Low-Rank Adaptation) (Hu et al., 2022; Houlsby et al., 2019) modifies the model’s attention and feedforward layers by injecting low-rank trainable matrices, further reducing the resources required for fine-tuning.
While Parameter-Efficient Fine-Tuning (PEFT) methods reduce training costs and memory usage, they may not always achieve the same level of task-specific performance as full model fine-tuning. Research has shown that for tasks requiring high accuracy, complex adaptations, or domain-specific knowledge, full fine-tuning often outperforms PEFT approaches due to its ability to adjust all model parameters for better adaptation (Hu et al., 2022; Li & Liang, 2021; Zaken et al., 2022). To make full model fine-tuning more computationally feasible, zeroth-order methods offer an alternative by reducing the high computational cost. Rather than computing gradients via backpropagation, zeroth-order methods estimate the gradient using only the forward pass. Initially explored in the 1990s (Spall, 1992; Nesterov & Spokoiny, 2017; Ghadimi & Lan, 2013; Duchi et al., 2015; Liu et al., 2020), these methods have recently gained traction for fine-tuning LLMs (Malladi et al., 2023; Deng et al., 2023a; Ling et al., 2024) and have been shown to be able to outperform first-order PEFT methods given enough training time (Zhang et al., 2024b).
However, zeroth-order methods often rely on simple, fixed prompts during fine-tuning. In tasks like sentiment analysis with the SST-2 dataset, templated prompts (e.g., $<CLS>$ text data. It was [terrible $\mid$ great]. $<SEP>$) are crucial for success (Zhang et al., 2024b). These prompts effectively align the text data with task-specific objectives. Therefore, selecting such templates becomes a key hyperparameter, raising the question: Can we automatically discover effective prompts for zeroth-order fine-tuning through prompt tuning? More broadly, can PEFT methods complement zeroth-order fine-tuning for large models? In this work, we propose a new framework to answer this question.
While our focus has thus far been on single-task fine-tuning, many scenarios necessitate multi-task fine-tuning. Multi-task learning (MTL) enables a model to handle multiple tasks simultaneously, fostering knowledge transfer between tasks and improving overall efficiency (Min et al., 2022; Yang et al., 2024). This approach is particularly valuable in low-resource settings, where collecting large labeled datasets can be costly, as is often the case with medical data. In such environments, few-shot learning—where a model is fine-tuned on a high-resource dataset to quickly adapt to new tasks with minimal data—becomes essential (Ye et al., 2021).
To address the challenges of multi-task and few-shot learning in natural language processing, several meta fine-tuning methods have been proposed (Huang et al., 2023; Zhao et al., 2024; Ye et al., 2021; Asadi et al., 2024). However, traditional meta fine-tuning approaches, such as MetaICL (Min et al., 2022), still require full-model first-order gradient calculations, which become computationally expensive with large language models (LLMs) containing billions of parameters. Given the success of zeroth-order methods in fine-tuning LLMs for individual tasks, the potential for adapting their applicability to multi-task learning remains largely unexplored.
With the effectiveness of zeroth-order fine-tuning and the advantages of PEFT for single tasks, a natural question arises: Can we develop a new multi-task and few-shot learning methodology that significantly reduces computational costs while maintaining or even enhancing performance? Specifically, can we leverage the efficiency of zeroth-order fine-tuning alongside the adaptability of PEFT within multi-task and few-shot learning for large language models?
1.1 Contributions
Bilevel Fine Tuning Framework: In this work we introduce a bilevel framework to enhance zeroth-order fine-tuning for large pre-trained language models using Parameter-Efficient Fine-Tuning (PEFT). It involves two levels of optimization: one for fine-tuning the base model and another for selecting the best PEFT parameters. The two levels are nested such that each level is enhancing the performance of the other level.
Efficient Training Method: We propose the Bilevel Zeroth-Order-First-Order (Bilevel ZOFO) method to solve the bilevel optimization problem. It incorporates zeroth-order approximations into first-order bilevel methods, avoiding full model gradient computation and addressing the high resource demands of traditional methods, especially for large models. We provide the convergence guarantees of the proposed method. Experiments demonstrate the superiority of Bilevel ZOFO, achieving better results than traditional PEFT and zeroth-order methods in single-task settings while maintaining similar memory efficiency.
Mitigating the sensitivity of hard prompts in ZO fine tuning The standard zeroth-order fine-tuning is very sensitive to the hard prompt used. Our method mitigates this sensitivity problem. In our experiments, the proposed method exhibits significantly lower sensitivity to the choice of hard prompts.
Light-weight Method for Multi-Task Learning: The framework is expanded to multi-task learning, where zeroth-order fine-tuning reduces computational costs, and PEFT models allow efficient task-specific fine-tuning. This results in a lightweight and adaptable meta-training process suitable for new tasks. In the experiments, our method outperforms existing approaches in multi-task learning scenarios.
2 Related work
2.1 Zeroth order in fine tuning LLMs
MeZO (Malladi et al., 2023) is the first work to use Zeroth-Order (ZO) methods to finetune LLMs for downstream tasks. They demonstrate that their method is compatible with both full-parameter tuning and parameter-efficient tuning techniques, such as LoRA and prefix tuning, while being significantly more computationally efficient. Zhang et al. (2024b) provide a benchmark for ZO optimization in the context of LLM fine-tuning, comparing different ZO optimizers and applying the method to various models. Gautam et al. (2024) introduce variance reduction techniques into ZO methods for fine-tuning, improving both stability and convergence. In addition, ZO methods are applied in federated fine-tuning by Qin et al. (2024) and Ling et al. (2024). Deng et al. (2023b) implement ZO optimization for softmax units in LLMs. Guo et al. (2024b) and Liu et al. (2024b) explore fine-tuning a minimal subset of LLM parameters using ZO methods by sparsifying gradient approximation or the perturbations used in gradient estimation. Tang et al. (2024) investigate the privacy of ZO optimization methods.
In contrast to previous approaches, we propose a bilevel training algorithm that effectively combines the strengths of both First-Order (FO) Parameter-Efficient Fine-Tuning (PEFT) and ZO full-model fine-tuning. Our experiments demonstrate that the bilevel structure, when paired with the most suitable PEFT technique, outperforms both ZO full-model fine-tuning and FO PEFT methods individually.
2.2 Fine-tuning LLMs for Multitask and Few-Shot Learning
Typical meta-tuning approaches employ First-Order methods to train autoregressive LLMs on a multitask dataset for various tasks (Zhong et al., 2021; Min et al., 2022; Guo et al., 2024a). Zhong et al. (2021) apply meta-training to tasks such as hate speech detection, question categorization, topic classification, and sentiment classification. Guo et al. (2024a) adopt the method from Min et al. (2022) for generating stylistic text. While Min et al. (2022) focus on enhancing the in-context learning ability of the meta-trained model for multitask learning, Zhong et al. (2021) focus on improving zero-shot performance.
During training, Min et al. (2022) sample a task from the dataset for each iteration to perform in-context learning. In contrast to Zhong et al. (2021) and Min et al. (2022), our approach uses a bilevel structure: the full LLM is fine-tuned at the upper level, while parameter-efficient fine-tuning (PEFT) models are tuned at the lower level. At test time, we freeze the meta-tuned base model and fine-tune only the PEFT model using a few-shot setup, which is both more cost-effective and efficient. Crucially, Min et al. (2022) fine tune the full model with first order methods, while we employ a ZO method in meta-tuning the base model at the upper level. Our approach allows us to bypass the need for backpropagation in the meta-model, significantly reducing computational costs.
2.3 Penalty Methods for Bilevel Optimization
Solving a bilevel optimization problem is challenging because the function value in the upper-level objective depends on the optimizer of the lower-level problem. This makes it difficult to compute the gradient of the upper-level objective, also known as the hypergradient. Classical methods require calculating Hessian-vector multiplications to approximate the hypergradient (Franceschi et al., 2017, 2018; Finn et al., 2017; Li et al., 2022; Rajeswaran et al., 2019; Ghadimi & Wang, 2018; Chen et al., 2022; Lorraine et al., 2020). However, when fine-tuning large language models, this process becomes extremely expensive due to the high computational and memory demands.
Recently, new frameworks for bilevel optimization have been introduced (Lu & Mei, 2024; Shen & Chen, 2023; Liu et al., 2024a; Kwon et al., 2023; Liu et al., 2022). These methods bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. The constraint is penalized, allowing the problem to be tackled as a minimax problem using only first-order information. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive.
Liu et al. (2024a) and Lu & Mei (2024) explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from Lu & Mei (2024) to approximate part of the upper-level parameters using a ZO approximation In order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
Figure 1: Bilevel ZOFO optimizes LLM fine-tuning by solving a bilevel problem using a penalty-based minimax approach, combining zeroth-order gradient estimation for LLM updates and first-order methods for PEFT parameters.
3 Bilevel model and Zeroth-order-first-order method
In this section, we present our bilevel model and the zeroth-order-first-order method for solving it. Let $\mathbf{p}$ represent the parameters of the PEFT model, and $\bm{\theta}$ represent the parameters of the pretrained base model. Given a single downstream task, such as classification, we aim to solve the following optimization problem:
$\min_{{\bm{\theta}}\in\mathbb{R}^{d}}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{f}),$ (1)
where ${\mathbf{p}}\in\mathbb{R}^{d^{\prime}}$ and $F({\bm{\theta}},{\mathbf{p}};\mathcal{B}):=\frac{1}{|\mathcal{B}|}\sum_{x\in\mathcal{B}}F({\bm{\theta}},{\mathbf{p}};x)$ is a loss function given a dataset $\mathcal{B}$.
When ${\mathbf{p}}$ corresponds to the embeddings of the hard prompt (as shown in Table 13 in the appendix of (Malladi et al., 2023)), the model above reduces to classical fine-tuning on a single downstream task. In model (1), the parameters of the PEFT model, ${\mathbf{p}}$, are fixed.
To enhance generalization ability, we split the dataset into two parts: one for tuning the PEFT model (denoted as $\mathfrak{D}_{\mathbf{p}}$) and another for fine-tuning the LLM (denoted as $\mathfrak{D}_{f}$). To maximize performance on downstream tasks, we need the optimal PEFT model parameters that are best suited for the current LLM base model. To achieve this, ${\mathbf{p}}$ should satisfy the following condition:
${\mathbf{p}(\bm{\theta})}\in\mathop{\rm arg\,min}_{{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}}).$
This condition reveals that as the parameters $\theta$ of the LLM change, the parameters ${\mathbf{p}}$ in the PEFT model should also be updated accordingly. Therefore, instead of solving (1), our true objective becomes:
$\begin{split}\min_{{\bm{\theta}}\in\mathbb{R}^{d}}&F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})\\ {\rm{\mathbf{s}}.t.\ }&{\mathbf{p}(\bm{\theta})}\in\mathop{\rm arg\,min}_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}}).\end{split}$ (2)
In this way, we find the optimal pair of parameters for both the PEFT model and the LLM base model to achieve the best performance on downstream tasks.
To solve the bilevel optimization problem (2), classical bilevel methods (as discussed in related work) view (2) as a single-level problem $\min_{\bm{\theta}}F({\bm{\theta}},{\mathbf{p}(\bm{\theta})})$. Since ${\mathbf{p}(\bm{\theta})}$ is the minimizer of another optimization problem, these methods typically require computing the Hessian-vector product (matrix multiplication of $\nabla_{{\bm{\theta}}{\mathbf{p}}}F({\bm{\theta}},{\mathbf{p}})$ and some vector $v$) multiple times to estimate the gradient of $F({\bm{\theta}},{\mathbf{p}(\mathbf{{\bm{\theta}}})})$ with respect to ${\bm{\theta}}$. However, for large language models (LLMs), this approach is computationally prohibitive because the number of parameters in ${\bm{\theta}}$ is too large.
To reduce the computational cost, following Lu & Mei (2024), we consider using a penalty method for the bilevel problem (2). Specifically, (2) is equivalent to the following constrained optimization problem:
$\begin{split}\min_{\bm{\theta}\in\mathbb{R}^{d},\bm{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}&F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{f})\\ {\rm{\mathbf{s}}.t.\ }&F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-\inf_{\mathbf{{\mathbf{s}}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}})\leq 0.\end{split}$ (3)
By penalizing the constraint, we obtain the following penalized problem:
$\min_{\begin{subarray}{c}\bm{\theta}\in\mathbb{R}^{d}\\ \bm{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}\end{subarray}}\!\!F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})+\lambda(F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-\!\!\inf_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}F({\bm{\theta}},{\mathbf{{\mathbf{s}}}};\mathfrak{D}_{\mathbf{p}})),$ (4)
where $\lambda>0$. As $\lambda$ increases, the solution to the penalized problem approaches the solution to (3), and thus the solution to (2) (see Lemma 5.4 for an explicit relationship between the stationary points of (4) and those of the original problem (2)). Note that the penalized problem (4) is equivalent to the following minimax problem:
$\min_{{\bm{\theta}}\in\mathbb{R}^{d},{\mathbf{p}}\in\mathbb{R}^{d^{\prime}}}\max_{{\mathbf{s}}\in\mathbb{R}^{d^{\prime}}}G_{\lambda}({\bm{\theta}},{\mathbf{p}},{\mathbf{s}}):=\\ F({\bm{\theta}},{\mathbf{p}(\bm{\theta})};\mathfrak{D}_{f})+\lambda\big{(}F({\bm{\theta}},{\mathbf{p}};\mathfrak{D}_{\mathbf{p}})-F({\bm{\theta}},{\mathbf{s}};\mathfrak{D}_{\mathbf{p}})\big{)}.$ (5)
In this way, we can solve the bilevel problem as a minimax problem. The basic minimax algorithm works as follows: at iteration $k$, we first solve the maximization problem $\max_{\mathbf{{\mathbf{s}}}}G_{\lambda}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}})$ with $({\bm{\theta}}^{k},{\mathbf{p}}^{k})$ fixed. For example, we can update ${\mathbf{s}}^{k}$ using an inner loop with stochastic gradient descent (SGD). Let ${\mathbf{s}}^{k+1}$ be the result of this inner loop. Then, in the outer loop, we update $({\bm{\theta}}^{k},{\mathbf{p}}^{k})$ by solving $\min_{{\bm{\theta}},{\mathbf{p}}}G_{\lambda}({\bm{\theta}},{\mathbf{p}},{\mathbf{s}}^{k+1})$ with ${\mathbf{s}}^{k+1}$ fixed. Again, SGD can be used to update ${\bm{\theta}}^{k}$ and ${\mathbf{p}}^{k}$. The conceptual algorithm is presented in Algorithm 1.
Algorithm 1 Bilevel first-order method
1:  Input: $\eta>0$, $\zeta>0$, ${\mathbf{s}}^{0}={\mathbf{s}}^{k}$, $K,T\in\mathbb{N}_{+}$, $\lambda\geq 0$.
2:  for k=0,…,K do
3:     for $t=0,\dots,T-1$do
4:        Let ${\mathbf{s}}^{k}_{t+1}={\mathbf{s}}^{k}_{t}-\eta\nabla_{\mathbf{{\mathbf{s}}}}G_{\lambda_{k}}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k}_{t})$.
5:        Output ${\mathbf{s}}^{k+1}={\mathbf{s}}^{k}_{T}$.
6:     end for
7:     Let ${\bm{\theta}}^{k+1}={\bm{\theta}}^{k}-\zeta\nabla_{\bm{\theta}}G_{\lambda_{k}}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k})$ and ${\mathbf{p}}^{k+1}={\mathbf{p}}^{k}-\zeta\nabla_{\mathbf{p}}G_{\lambda}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k}).$
8:  end for
However, note that
$\nabla_{\bm{\theta}}G_{\lambda_{k}}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k})=\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})+\\ \lambda_{k}(\nabla_{{\bm{\theta}}}F({\bm{\theta}},{\mathbf{p}}^{k};\mathfrak{D}_{\mathbf{p}})+\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k};\mathfrak{D}_{\mathbf{p}})),$ (6)
requires calculating the gradient with respect to ${\bm{\theta}}$, i.e, $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})$. Given the large scale of ${\bm{\theta}}$ in LLMs, this is computationally expensive. To avoid this, we use zeroth-order (ZO) information to approximate the gradient $\nabla_{\bm{\theta}}G$. Following (Malladi et al., 2023; Zhang et al., 2024b; Guo et al., 2024b), we employ the Simultaneous Perturbation Stochastic Approximation (SPSA) as a classical zeroth-order gradient estimator. Specifically, at each iteration $k$, we sample ${\mathbf{z}}^{k}\sim N(0,I_{d})$, where $d$ is the dimension of ${\bm{\theta}}$. We then approximate the gradient $\nabla_{\bm{\theta}}F$ as follows:
^$\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};x):=\\ \frac{F({\bm{\theta}}^{k}+\epsilon{\mathbf{z}}^{k},{\mathbf{p}}^{k};x)-F({\bm{\theta}}^{k}-\epsilon{\mathbf{z}}^{k},{\mathbf{p}}^{k};x)}{2\epsilon}{\mathbf{z}}^{k}.$ (7)
As opposed to the number of LLM parameters ${\bm{\theta}}$, the number of PEFT parameters ${\mathbf{p}}$ is very small. So it is feasible to compute the exact gradient with respect to ${\mathbf{p}}$. Thus, we calculate $\nabla_{\mathbf{p}}F({\bm{\theta}},{\mathbf{p}};\mathcal{B})$ exactly.
Additionally, in each iteration $k$ and inner iteration $t$ of Algorithm 1, we sample a mini-batch $\mathcal{B}$ of size $B$. We use ^$\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B})$ to substitude $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{D}_{f})$ and $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{D}_{p})$in (6). We also use mini-batches when calculating the gradients with respect to the PEFT parameters ${\mathbf{s}}$ and ${\mathbf{p}}$.
This approach leads to the final algorithm (Algorithm 2 and Figure 1) for fine-tuning LLMs using the bilevel model (2). We refer to this method as the Bilevel Zeroth-Order-First-Order (Bilevel ZOFO) method.
Algorithm 2 Bilevel Zeroth-order-first-order Method (Bilevel ZOFO)
1:  Input: $\eta>0$, $\zeta>0$, batchsize $B$, ${\mathbf{s}}^{0}={\mathbf{s}}^{k}$, $K,T\in\mathbb{N}_{+}$, $\lambda>0$.
2:  for k=0,…,K do
3:     for t=0,…,T-1 do
4:        Sample a batch $\mathcal{B}^{k}_{t,{\mathbf{p}}}$ from $\mathfrak{D}_{\mathbf{p}}$.
5:        Let ${\mathbf{s}}^{k}_{t+1}={\mathbf{s}}^{k}_{t}-\eta\nabla_{\mathbf{{\mathbf{s}}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k}_{t};\mathcal{B}^{k}_{t,{\mathbf{p}}})$
6:        Output ${\mathbf{s}}^{k+1}={\mathbf{s}}^{k}_{T}$.
7:     end for
8:     Sample a batch $\{\mathcal{B}^{k}_{f}\}$ from $\mathfrak{D}_{f}$ and $\{\mathcal{B}^{k}_{\mathbf{p}}\}$ from $\mathfrak{D}_{\mathbf{p}}$.
9:     For $x\in\mathcal{B}^{k}_{\mathbf{p}}\cup\mathcal{B}^{k}_{f}$, calculate ^$\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};x)$ following (7).
10:     Let
${\mathbf{p}}^{k+1}={\mathbf{p}}^{k}-\zeta(\nabla_{{\mathbf{p}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{f})+\\ \lambda_{k}(\nabla_{{\mathbf{p}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{\mathbf{p}})))$ (8)
${\bm{\theta}}^{k+1}={\bm{\theta}}^{k}-\zeta(\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{f})+\\ \lambda_{k}(\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{\mathbf{p}})-\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k+1};\mathcal{B}^{k}_{\mathbf{p}})))$ (9)
11:  end for
4 Experiments
We conduct extensive experiments on various LLMs of different scales to demonstrate the effectiveness of bilevel-ZOFO in both single-task and multi-task meta-training settings. Our proposed structure is able to incorporate any variation of zeroth-order methods in the upper-level step and any PEFT method in the lower level. To maintain focus on testing the effectiveness of the proposed bilevel structure and its unique multitask learning capabilities, we used the classic MeZO  (Malladi et al., 2023).
4.1 Bilevel-ZOFO for Single-task Fine-tuning
Following MeZO (Malladi et al., 2023), we evaluate our approach on a range of classification and multiple-choice tasks. In this setting, training and testing are conducted on the same task. We employ prompt-tuning (Lester et al., 2021), prefix-tuning (Li & Liang, 2021), and LoRA (Hu et al., 2022)- well-known PEFT baselines-for lower-level training to validate bilevel-ZOFO under different conditions and resource constraints. During each lower-level update, we update only the PEFT parameters, and during the upper-level optimization step, we tune the full model using zeroth-order gradient approximation. We perform 10 lower-level updates between each pair of upper-level updates. For each task, we randomly sample 1000 examples for training, 500 examples for validation, and 1000 examples for testing. We use the Adam optimizer (Kingma & Ba, 2015) and report test accuracy or F1-score.
We compare our method against several baselines, including MeZO for Full Model Fine-tuning, MeZO for PEFT, and First-order PEFT. We fix the total memory budget of each step across bilevel-ZOFO and the baselines. We train zeroth-order methods for 10,000 steps, and first-order methods for 5000 steps. For all experimental details, refer to the Appendix B.1.3 and Appendix B.1.4. We also provide the training loss for the lower-level objective of the bilevel framework in Figure 4 of Appendix B.2 to show that consistent with the guarantees provided by our theoretical analysis in Section 5, Bilevel-ZOFO converges. We compare the memory requirements of our method with the baselines in Figure 5, and provide wall-clock analysis in Table 7. Figure 5 and Table 7 show that bilevel-ZOFO does not increase the memory usage of PEFT and MeZO.
Bileve-ZOFO mitigates sensitivity to hard prompts: We begin our experiments by showing that Bilevel-ZOFO mitigates the sensitivity of MeZO to hard prompts. First, similar to Table 5 of the MeZO (Malladi et al., 2023) paper, we experiment with different choices of hard prompts for both MeZO and our Bilevel-ZOFO. Table 1 shows the results for tuning Opt 1.3b on SST2 and COPA. We can see that our method effectively mitigates the sensitivity of MeZO to hard prompts. The difference between the results with and without a simple hard prompt in our experiment is much less than MeZO’s. See Appendix B.4 for a full discussion.
Method Task w/ prompt (%) w/o prompt (%) Diff.
MeZO SST-2 89.6 51.9 -38.6
COPA 70.0 54.8 -15.2
Bilevel-ZOFO SST-2 93.3 92.9 -0.4
COPA 76.66 73.6 -3.06
Table 1: Prompt Sensitivity Comparison for MeZO and Bilevel-ZOFO. Bilevel-ZOFO effectively mitigates the sensitivity of MeZO to hard prompts.
Trainer Mode BoolQ CB Copa ReCoRD RTE SST2 WIC WinoGrande WSC Average
MeZO ft 0.6927 0.7767 0.7000 0.6980 0.6587 0.8214 0.5543 0.5480 0.5054 0.6617
lora 0.6860 0.7607 0.7200 0.7083 0.6755 0.8501 0.5549 0.5607 0.5570 0.6748
prefix 0.6573 0.7945 0.7033 0.7047 0.6972 0.8218 0.5622 0.5370 0.5105 0.6654
prompt 0.6260 0.5821 0.7067 0.7070 0.5415 0.7463 0.5574 0.5556 0.4654 0.6098
average 0.6655 0.7285 0.7075 0.7045 0.6432 0.8099 0.5572 0.5503 0.5096 0.6529
FO lora 0.7456 0.8512 0.7500 0.7206 0.7292 0.9258 0.6463 0.5806 0.6474 0.7330
prefix 0.7300 0.8571 0.7167 0.7093 0.7136 0.8133 0.5387 0.5787 0.5705 0.6920
prompt 0.7150 0.7142 0.7466 0.7163 0.6936 0.8016 0.5386 0.5980 0.5062 0.6700
average 0.7302 0.8075 0.7378 0.7154 0.7121 0.8470 0.5745 0.5857 0.5747 0.6977
[…]
Table 2: Single-Task Experiments on OPT-1.3B with 1000 samples. Values correspond to mean across three random seeds. FO: First-Order. FT: full-model fine-tuning. See Table 5 in the Appendix for standard deviation values.
Table 2 presents the test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B (Zhang et al., 2022) on a downstream task. Table 3 demonstrates the results for Llama2-7b (Touvron et al., 2023). We can make the following observations:
Bilevel-ZOFO outperforms MeZO on almost all tasks: With the same memory allocation per training step, bilevel-ZOFO outperforms MeZO, even when trained for half the number of iterations across most tasks.
Bilevel-ZOFO outperforms FO PEFT on average:
Comparing each FO-PEFT setting with the corresponding bilevel-ZOFO setting, we see that bilevel-ZOFO outperforms the corresponding FO-PEFT methods across most instances and on average.
Bilevel-ZOFO outperforms baselines more significantly in resource-constrained settings: The number of parameters tuned for prefix tuning and prompt tuning is lower than for LoRA (See Figure 3). As shown in Table 2, when fewer parameters are tuned, bilevel-ZOFO demonstrates a larger improvement over first-order methods in tuning PEFT models. Since memory usage and training steps remain the same, bilevel-ZOFO proves to be a more suitable option for fine-tuning LLMs in constrained environments compared to PEFT and MeZO.
Bilevel-ZOFO scales effectively to larger LLMs: Table 3 compares bilevel-ZOFO with the baselines when fine-tuning Llama2-7b (Touvron et al., 2023) on various classification and open-ended generation tasks. The results show that bilevel-ZOFO’s advantages are not confined to smaller models like OPT-1.3b, but also extend to larger LLMs.
4.2 Multi-Task Fine-Tuning Experiments
Following Min et al. (2022), we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings: classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. For instance, in non-classification-to-classification setting, we train on a number of non-classification subtasks and test on a number of distinct classification subtasks. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT (Ye et al., 2021) and UNIFIEDQA (Khashabi et al., 2020), comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification and question answering all in English. We use GPT2-Large (Radford et al., 2019) as the base model for these experiments.
We compare our method against several baseline approaches:
  • MetaICL (Min et al., 2022): A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, ${(a_{1},b_{1}),\dots,(a_{k},b_{k})}$, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
  • Zero-shot: This method uses the pretrained language model (LM) without any tuning, performing zero-shot inference without any demonstration examples.
  • In-context Learning (ICL): This method uses the pretrained LM with in-context learning by conditioning on a concatenation of $k$ demonstration examples and 1 actual test sample similar to MetaICL.
We sample 768 examples from each training sub-task. We train MetaICL in their original setting for 30,000 steps. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we DO NOT include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  (Malladi et al., 2023), we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on 4 demonstration samples from each class of each test sub-task. Similar to (Min et al., 2022), we report Macro-averaged F1 as the evaluation metric. See Appendix B.5 for all training details.
Table 4 presents the meta-learning results. We observe that in zero-shot setting, bilevel-ZOFO (ours(zeroshot)) outperforms zero-shot on all tasks. Note that although ICL and MetaICL perform better than ours (zero-shot) 1)MetaICL fine-tunes the entire base model using first-order methods, which incurs a significantly higher computational cost. 2)both ICL and MetaICL with $k=4$ demonstration examples take 4 times more time to do inference than our method with no demonstration examples. Nonetheless, after a lightweight 10-iteration LoRA fine-tuning phase, bilevel-ZOFO(ours(tuned)) surpasses ICL and MetaICL on nearly every hyper-task, highlighting its strong potential as a meta-learning algorithm.
Trainer Mode BoolQ ReCoRD SQuAD SST2 Average
MeZO ft 0.7915 0.7890 0.7737 0.8646 0.8047
lora 0.8020 0.7970 0.7412 0.8529 0.7983
prefix 0.7830 0.7905 0.7093 0.8364 0.7798
prompt 0.7787 0.7935 0.7014 0.8246 0.7746
average 0.7888 0.7925 0.7489 0.8397 0.7825
FO lora 0.8420 0.7920 0.8197 0.9557 0.8524
prefix 0.7783 0.8013 0.7946 0.9243 0.8246
prompt 0.8083 0.8023 0.7805 0.9284 0.8299
average 0.8095 0.7985 0.7983 0.9361 0.8356
[…]
Table 3: Single-Task Experiments on Llama2-7B with 1000 samples. Values correspond to mean across three random seeds. FO: First-Order. FT: full-model fine-tuning. See Table 6 for full details.
Method class non_class qa non_qa
$\rightarrow$ class $\rightarrow$ class $\rightarrow$ qa $\rightarrow$ qa
Zero-shot 34.2 34.2 40.2 40.2
Few-shot 34.9 (1.4) 34.9 (1.4) 40.5 (0.3) 40.5 (0.4)
MetaICL 46.4 (1.1) 37.7 (1.7) 45.5 (0.3) 40.2 (0.6)
Ours (Zero-shot) 34.5 34.3 41.8 40.4
Ours(Tuned) 47.1 42.4 43.5 (1.3) 41.9
Table 4: Multi-task Meta learning results using GPT2-Large as the base model. Values correspond to the mean and standard deviation over 5 test seeds which include different demonstration samples for each test task. class: Classification, qa: Question Answering
4.3 Ablative Studies
We perform an ablation study by varying the regularization parameter $\lambda$ (as defined in Equation (5)) and the number of lower-level training steps between each pair of upper-level updates. Figure 2 shows the results. From Figure 2(a), the effect $\lambda$ appears to be non-linear, indicating the need to find an optimal balance. Nontheless, a moderate value like $10$ or $100$ seems to work reasonably well on all tasks. As anticipated, Figure 2(b) demonstrates that performance generally degrades when the total number of upper-level updates is reduced, suggesting there is a trade-off between latency and performance. While more upper-level updates improve results, they also extend the overall training time.
(a) $\lambda$
(b) Number of lower-level training steps
Figure 2: Ablation over $\lambda$ in (5) and the number of lower-level training steps before each upper-level update.
5 Analysis
In this section we give convergence guarantee for Bilevel ZOFO. Suppose $({\mathbf{{\bm{\theta}}}},{\mathbf{p}})\in\mathbb{R}^{d+d^{\prime}}$ and ${\mathbf{s}}\in\mathbb{R}^{d^{\prime}}$. The following assumptions are made throughout this section.
Assumption 5.1.
We make the following assumptions:
  • $G({\mathbf{{\bm{\theta}}}},{\mathbf{{\mathbf{p}}}},\cdot)$ can be potentially nonconvex and $G(\cdot,\cdot,{\mathbf{s}})$ is $\tau-$ strongly concave; $F({\mathbf{{\bm{\theta}}}},{\mathbf{p}})$ is twice continuously differentiable in ${\mathbf{{\bm{\theta}}}},{\mathbf{p}}$.
  • $G$ is $\ell$-Lipschitz smooth in $\mathbb{R}^{d+2d^{\prime}},$ i.e. $\forall({\mathbf{{\bm{\theta}}}}_{1},{\mathbf{p}}_{1},{\mathbf{s}}_{1}),({\mathbf{{\bm{\theta}}}}_{2},{\mathbf{p}}_{2},{\mathbf{s}}_{2})\in\mathbb{R}^{d+2d^{\prime}}$,
    $\|\nabla G({\mathbf{{\bm{\theta}}}}_{1},{\mathbf{p}}_{1},{\mathbf{s}}_{1})-\nabla G({\mathbf{{\bm{\theta}}}}_{2},{\mathbf{p}}_{2},{\mathbf{s}}_{2})\|\leq\\ \ell\|({\mathbf{{\bm{\theta}}}}_{1},{\mathbf{p}}_{1},{\mathbf{s}}_{1})-({\mathbf{{\bm{\theta}}}}_{2},{\mathbf{p}}_{2},{\mathbf{s}}_{2})\|.$
    We define $\kappa:=\ell/\tau$ as the problem condition number.
  • $\forall({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})\in\mathbb{R}^{d+2d^{\prime}}$, sample estimates satisfy
    $\begin{split}&\mathbb{E}[G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}};\xi)]=G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}}),\\ &\mathbb{E}[\nabla G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}};\xi)]=\nabla G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}}),\\ &\mathbb{E}\|\nabla G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}};\xi)-\nabla G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})\|^{2}\leq\frac{\sigma^{2}}{B}\end{split}$
    for sample $\xi$ with size $|\xi|=B$ and constant $\sigma>0$.
  • $\max_{\mathbf{s}}G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})$ is lower bounded.
We first discuss the relationship between the optimality condition (4) and (2). We start with defining the $\epsilon$-stationary points of (4) and (2) for general bilevel and minimax problems. In the following definitions, the expectation is taken over the randomness in the algorithm that $({\mathbf{x}},{\mathbf{y}})$ is generated.
Definition 5.2.
Given a bilevel optimization problem
$f^{*}=\min_{\mathbf{x}}f({\mathbf{x}},{\mathbf{y}}^{*}({\mathbf{x}})),{\mathbf{y}}^{*}({\mathbf{x}})\in\arg\min_{\mathbf{z}}g({\mathbf{x}},{\mathbf{z}})$
and any $\epsilon>0$, a point $({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})$ is called an $\epsilon$-stationary point if
$\mathbb{E}[\|\nabla f({\mathbf{x}}_{\epsilon},{\mathbf{y}}^{*}({\mathbf{x}}_{\epsilon}))\|]\leq O(\epsilon),f({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})-\min_{\mathbf{z}}f({\mathbf{x}}_{\epsilon},{\mathbf{z}})\leq\epsilon.$
Definition 5.3.
Given a minimax problem
$f^{*}=\min_{\mathbf{x}}\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$
and any $\epsilon>0$, a point $({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})$ is called an $\epsilon$-stationary point if
$\mathbb{E}[\|\nabla_{\mathbf{x}}f({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})\|^{2}]\leq\epsilon^{2},\ \mathbb{E}[\|\nabla_{\mathbf{y}}f({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})\|^{2}]\leq\epsilon^{2}.$
Lemma 5.4.
If assumption 5.1 holds and $\lambda=1/\epsilon$, assume that $\nabla^{2}F({\mathbf{{\bm{\theta}}}},\cdot)$ is Lipschitz continuous and $({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})$ is an $\epsilon$-stationary point of (4), then $({\mathbf{{\bm{\theta}}}},{\mathbf{s}})$ is an $\epsilon$-stationary point of (2).
The following is the low effective rank assumption from (Malladi et al., 2023). This assumption avoids dimension $d$ in the total complexity. Following (Malladi et al., 2023), we assume here that ${\mathbf{z}}^{k}$ in (7) is sampled from shpere in $\mathbb{R}^{d}$ with radius $\sqrt{d}$ for ease of illustration.
Assumption 5.5.
For any $({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})\in\mathbb{R}^{d+2d^{\prime}}$, there exists a matrix $H({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})$ such that $\nabla^{2}G({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})\preceq H({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})\preceq\ell\cdot I_{d}$ and $tr(H({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}}))\leq r\cdot\|H({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})\|.$
Theorem 5.6.
If Assumptions 5.1 and 5.5 hold, by setting
$\begin{split}&\eta=\frac{1}{2\ell},\zeta=\frac{1}{2\ell r},\lambda=\frac{1}{\epsilon},B=O(\sigma^{2}\epsilon^{-2}),\\ &\alpha=O(\epsilon\kappa^{-1}(d+d^{\prime})^{-1.5}),T=O\left(\kappa\log(\kappa\epsilon^{-1})\right),\\ &K=O(\kappa r\epsilon^{-2})\end{split}$
there exists an iteration in Algorithm 2 that returns an $\epsilon$-stationary point $({\mathbf{{\bm{\theta}}}},{\mathbf{p}},{\mathbf{s}})$ for (5) and it satisfies
$\begin{split}&\mathbb{E}[\|\nabla F({\mathbf{{\bm{\theta}}}},{\mathbf{p}}^{*}({\mathbf{{\bm{\theta}}}});\mathcal{D}_{f})\|]\leq O(\epsilon),\\ &F({\mathbf{{\bm{\theta}}}},{\mathbf{s}};\mathcal{D}_{\mathbf{p}})-\min_{\mathbf{p}}F({\mathbf{{\bm{\theta}}}},{\mathbf{p}};\mathcal{D}_{\mathbf{p}})\leq\epsilon.\end{split}$
Remark 5.7.
The total number of ZO gradient calculations is
$TKB_{1}+KB_{2}=O(\sigma^{2}\kappa^{2}r\epsilon^{-4}\log(\kappa\epsilon^{-1})).$
This result matches the complexity in previous ZO minimax algorithm in (Wang et al., 2023) but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption 5.5, providing efficiency guarantee for our algorithm.
6 Conclusions
In this work, we introduced a novel bilevel optimization framework designed to mitigate the downsides of PEFT and zeroth-order full model fine-tuning. We propose a new method that is more efficient than existing bilevel methods and thus more suitable for tuning full pre-trained large language models. Theoretically, we provide convergence guarantees for this new method. Empirically, we show that this method outperforms both zeroth-order and FO PEFT methods in single task settings. Additionally, we show this method is effective and efficient when adapted to do multi-task learning. With competitive and even better performance compared to existing meta-training methods, our method offers a significantly cheaper training process.
Impact Statement
This paper introduces an optimization framework that enhances the efficiency of fine-tuning large language models. By reducing computational costs and memory requirements, the approach promotes environmental sustainability and broadens access to advanced AI tools, promoting accessibility in AI development.
Appendix A Method
A.1 Proofs
In the proofs we use the simplified notations ${\mathbf{x}}:=({\bm{\theta}},{\mathbf{p}})$, ${\mathbf{y}}:={\mathbf{s}}$, $f({\mathbf{x}},{\mathbf{y}}):=G({\bm{\theta}},{\mathbf{p}},{\mathbf{s}})$, ${\mathbf{y}}^{*}({\mathbf{x}}):=\arg\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$ and $g({\mathbf{x}}):=f({\mathbf{x}},{\mathbf{y}}^{*}({\mathbf{x}})).$
A.1.1 proof of lemma 5.4
First we introduce some lemmas from previous literature.
Lemma A.1.
(Lemma 1.2.3, Theorem 2.1.8 and Theorem 2.1.10 in (Nesterov, 2013))
  • Suppose a function $h$ is $L_{h}$-gradient-Lipschitz and has a unique maximizer ${\mathbf{x}}^{*}$. Then, for any ${\mathbf{x}}$, we have:
    $\frac{1}{2L_{h}}\|\nabla h({\mathbf{x}})\|_{2}^{2}\leq h({\mathbf{x}}^{*})-h({\mathbf{x}})\leq\frac{L_{h}}{2}\|{\mathbf{x}}-{\mathbf{x}}^{*}\|_{2}^{2}.$
  • Suppose a function $h$ is $\tau_{h}$-strongly concave and has a unique maximizer ${\mathbf{x}}^{*}$. Then, for any ${\mathbf{x}}$, we have:
    $\frac{\tau_{h}}{2}\|{\mathbf{x}}-{\mathbf{x}}^{*}\|_{2}^{2}\leq h({\mathbf{x}}^{*})-h({\mathbf{x}})\leq\frac{1}{2\tau_{h}}\|\nabla h({\mathbf{x}})\|_{2}^{2}.$
From lemma A.1 and the definition of $\epsilon$-stationary point (in definition 5.3) we can get the following lemma.
Lemma A.2.
Suppose assumption 5.1 holds and $({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})$ is an $\epsilon$-stationary point of $\min_{\mathbf{x}}\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$, let $({\bm{\theta}}_{\epsilon},{\mathbf{p}}_{\epsilon})={\mathbf{x}}_{\epsilon}$ we have
$F({\bm{\theta}}_{\epsilon},{\mathbf{s}}_{\epsilon})-\min_{\mathbf{s}}F({\bm{\theta}}_{\epsilon},{\mathbf{s}})\leq O(\frac{\epsilon^{2}}{\lambda^{2}}).$
Proof.
$F({\bm{\theta}}_{\epsilon},{\mathbf{s}}_{\epsilon})-\min_{\mathbf{s}}F({\bm{\theta}}_{\epsilon},{\mathbf{s}})\leq\frac{1}{\tau}\|\nabla_{\mathbf{s}}F({\bm{\theta}}_{\epsilon},{\mathbf{s}}_{\epsilon})\|^{2}=\frac{1}{\lambda^{2}\tau}\|\nabla_{\mathbf{y}}f({\mathbf{x}}_{\epsilon},{\mathbf{y}}_{\epsilon})\|^{2}\leq O(\frac{\epsilon^{2}}{\lambda^{2}}),$
here the first inequality is from Lemma A.1 applied to $-F$ and the second inequality from definition 5.3. ∎
The following is a rephrase of theorem 2 in (Lu & Mei, 2024).
Proof.
(proof of lemma 5.4) By Lemma A.2 and the value of $\lambda$ we have
$F({\bm{\theta}}_{\epsilon},{\mathbf{s}}_{\epsilon})-\min_{\mathbf{s}}F({\bm{\theta}}_{\epsilon},{\mathbf{s}})\leq O(\epsilon^{4}).$
Therefore, by Theorem 2 in (Lu & Mei, 2024) we have $\mathbb{E}[\|\nabla F({\bm{\theta}},{\mathbf{p}}^{*}({\bm{\theta}}))\|]\leq O(\epsilon)$ and Lemma 5.4 is proven. ∎
A.1.2 proof of theorem 5.6
Based on Lemma 5.4, it suffices to prove that the algorithm 2 outputs an $\epsilon$-stationary point of $\min_{\mathbf{x}}\max_{\mathbf{y}}f({\mathbf{x}},{\mathbf{y}})$. In this section we will prove this conclusion.
First we introduce the smoothed function of $f$, which will be useful in the proof.
Lemma A.3.
(Lemma C.2 in (Zhang et al., 2024a)) Let ${\mathbf{u}}$ be uniformly sampled from the Euclidean sphere $\sqrt{d}\mathbb{{\mathbf{s}}}^{d-1}$ and ${\mathbf{v}}$ be uniformly sampled from the Euclidean ball $\sqrt{d}\mathbb{B}^{d}=\{{\mathbf{x}}\in\mathbb{R}^{d}\mid\|{\mathbf{x}}\|\leq\sqrt{d}\}$. For any function $f({\mathbf{x}}):\mathbb{R}^{d}\to\mathbb{R}$ and $\alpha>0$, we define its zeroth-order gradient estimator as:
^$\hat{\nabla}f_{\alpha}({\mathbf{x}})=\frac{f({\mathbf{x}}+\alpha{\mathbf{u}})-f({\mathbf{x}}-\alpha{\mathbf{u}})}{2\alpha}{\mathbf{u}},$
and the smoothed function as:
$f_{\alpha}({\mathbf{x}})=\mathbb{E}_{\mathbf{v}}[f({\mathbf{x}}+\alpha{\mathbf{v}})].$
The following properties hold:
  • (i)
    $f_{\alpha}({\mathbf{x}})$ is differentiable and $\mathbb{E}_{\mathbf{u}}[\hat{\nabla}f_{\alpha}({\mathbf{x}})]=\nabla f_{\alpha}({\mathbf{x}})$.
  • (ii)
    If $f({\mathbf{x}})$ is $\ell$-smooth, then we have that:
    $\|\nabla f({\mathbf{x}})-\nabla f_{\alpha}({\mathbf{x}})\|\leq\frac{\ell}{2}\alpha d^{3/2}.$
If we use $f({\mathbf{x}},{\mathbf{y}};\xi)$ to denote a forward evaluation with random samples $\xi$ and let batch size $B=|\xi|$, then $f({\mathbf{x}},\cdot;\xi)$ is a function from $\mathbb{R}^{d}$ to $\mathbb{R}$ and $\ell$-smooth. The above lemma can be used on $f({\mathbf{x}},\cdot)$ and $f({\mathbf{x}},\cdot;\xi)$. We can define its smoothed function $f_{\alpha}({\mathbf{x}},\cdot;\xi)$ and has the properties above.
Lemma A.4.
If assumption 5.1 holds, for $f_{\alpha}$ defined in Lemma A.3, $\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}},{\mathbf{y}})$ is $\ell$-continuous on ${\mathbf{y}}$, i.e.
$\|\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}},{\mathbf{y}}_{1})-\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}},{\mathbf{y}}_{2})\|\leq\ell\|{\mathbf{y}}_{1}-{\mathbf{y}}_{2}\|,$
for any ${\mathbf{x}}\in\mathbb{R}^{d},{\mathbf{y}}_{1},{\mathbf{y}}_{2}\in\mathbb{R}^{d^{\prime}}$.
Proof.
$\|\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}},{\mathbf{y}}_{1})-\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}},{\mathbf{y}}_{2})\|$
$=$ $\|\mathbb{E}_{\mathbf{v}}[f({\mathbf{x}}+\alpha{\mathbf{v}},{\mathbf{y}}_{1})]-\mathbb{E}_{\mathbf{v}}[f({\mathbf{x}}+\alpha{\mathbf{v}},{\mathbf{y}}_{2})]\|$
$\leq$ $\mathbb{E}_{\mathbf{v}}\|f({\mathbf{x}}+\alpha{\mathbf{v}},{\mathbf{y}}_{1})-f({\mathbf{x}}+\alpha{\mathbf{v}},{\mathbf{y}}_{2})\|$
$\leq$ $\ell\|{\mathbf{y}}_{1}-{\mathbf{y}}_{2}\|.$
Here the first inequality is from the convexity of norm and the second inequality is from the $\ell$-smoothness of $f$. ∎
We first give the iteration complexity of the inner loop of Algorithm 2. Using the simplified notations we can write the update step in the inner loop as ${\mathbf{y}}^{k}_{t+1}={\mathbf{y}}^{k}_{t}+\eta\nabla_{\mathbf{y}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t};\xi_{t}).$ We use $B_{1},B_{2}$ to denote the batch size for the inner loop and outer loop, respectively. But finally we will prove that they are in fact of the same order.
Lemma A.5.
In Algorithm 2, by setting $\eta=1/2\ell$, $T=O(\kappa\log(\frac{1}{\epsilon}))$ and $B_{1}=O(\epsilon^{-2})$ we have
$\mathbb{E}[\|{\mathbf{y}}^{k}_{T}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]\leq\epsilon^{2}$
in outer loop $k$.
Proof.
$\|{\mathbf{y}}^{k}_{t+1}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}$
$=$ $\|{\mathbf{y}}^{k}_{t}+\eta\nabla_{\mathbf{y}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t};\xi_{t})-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}$
$=$ $\|{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}+2\eta\langle\nabla_{\mathbf{y}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t};\xi_{t}),{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\rangle+\eta^{2}\|\nabla_{\mathbf{y}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t};\xi_{t})\|^{2}.$
Now taking expectations on both sides we have
$\mathbb{E}[\|{\mathbf{y}}^{k}_{t+1}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]$
$\leq$ $\mathbb{E}[\|{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]+2\eta\mathbb{E}[\langle\nabla_{\mathbf{y}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t}),{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\rangle]+\eta^{2}(\mathbb{E}[\|\nabla_{\mathbf{y}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t})\|^{2}]+\frac{\sigma^{2}}{B_{1}})$
$\leq$ $\mathbb{E}[\|{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]-2\eta\mathbb{E}[f({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))-f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t})]+2\ell\eta^{2}\mathbb{E}[f({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))-f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t})]+\frac{\eta^{2}\sigma^{2}}{B_{1}}$
$=$ $\mathbb{E}[\|{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]-\frac{1}{2\ell}\mathbb{E}[f({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))-f({\mathbf{x}}^{k},{\mathbf{y}}^{k}_{t})]+\frac{\sigma^{2}}{4\ell^{2}B_{1}}$
$\leq$ $\mathbb{E}[\|{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]-\frac{\tau}{4\ell}\mathbb{E}[\|{\mathbf{y}}^{k}_{t}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]+\frac{\sigma^{2}}{4\ell^{2}B_{1}}.$
The first inequality is from Assumption 5.1, second and last inequalities from Lemma A.1 and the equation is from the value of $\eta$.
In order for $\mathbb{E}[\|{\mathbf{y}}^{k}_{T}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]\leq\epsilon^{2}$ we need $T=O(\kappa\log(\frac{1}{\epsilon}))$ and $B_{1}=O(\epsilon^{-2}).$
The following lemma is from Theorem 1 in (Malladi et al., 2023).
Lemma A.6.
If Assumption 5.5 holds, there exists a constant $\gamma={\bm{\theta}}(r)$ such that
$\mathbb{E}[\hat{\nabla}_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)^{T}H({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\hat{\nabla}_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)]\leq\ell\gamma\mathbb{E}[\|\nabla_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)\|^{2}].$
Finally, we give the proof for Theorem 5.6. In this part we assume both ${\bm{\theta}}$ and ${\mathbf{p}}$ updates with zeroth order gradient for the convenience of analysis and this does not change the order of the total complexity.
Proof.
(proof of Theorem 5.6)
From Assumption 5.5, taking expectation conditioning on ${\mathbf{x}}^{k}$ and ${\mathbf{y}}^{k+1}$ we have
$\mathbb{E}[g({\mathbf{x}}^{k+1})]\leq$ $g({\mathbf{x}}^{k})-\zeta\langle\nabla_{\mathbf{x}}g({\mathbf{x}}^{k}),\mathbb{E}[\hat{\nabla}_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)]\rangle$
$+\frac{\zeta^{2}}{2}\mathbb{E}[\hat{\nabla}_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)^{T}H({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\hat{\nabla}_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)]$
$\leq$ $g({\mathbf{x}}^{k})-\zeta\langle\nabla_{\mathbf{x}}g({\mathbf{x}}^{k}),\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\rangle+\frac{\zeta^{2}}{2}\ell\gamma\mathbb{E}[\|\nabla_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)\|^{2}]$
Let us bound the inner product term:
$-\zeta\langle\nabla_{\mathbf{x}}g({\mathbf{x}}^{k}),\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\rangle$
$\leq$ $-\zeta\langle\nabla_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))-\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))+\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))$
$-\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})+\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1}),\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\rangle$
$\leq$ $\frac{1}{\ell\gamma}\|\nabla_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))-\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))\|^{2}+\frac{\zeta^{2}\ell\gamma}{4}\|\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}$
$+\frac{1}{\ell\gamma}\|\nabla_{\mathbf{x}}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k}))-\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}+\frac{\zeta^{2}\ell\gamma}{4}\|\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}$
$-\zeta\langle\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1}),\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\rangle$
$\leq$ $\frac{\alpha^{2}\ell^{2}d^{3}}{4\ell\gamma}+\frac{\ell^{2}}{\ell\gamma}\|{\mathbf{y}}^{*}({\mathbf{x}}^{k})-{\mathbf{y}}^{k+1}\|^{2}+\frac{\zeta^{2}\ell\gamma}{2}\|\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}$
$-\zeta\langle\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1}),\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\rangle.$
Here the last inequality is from Lemma A.3 and Lemma A.4.
Now back to the original inequality, taking expectations over all the randomness in the algorithm we have
$\zeta(1-\frac{\zeta\ell\gamma}{2})\mathbb{E}[\|\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]$
$\leq$ $\mathbb{E}[g({\mathbf{x}}^{k})-g({\mathbf{x}}^{k+1})]+\frac{\ell}{\gamma}\mathbb{E}[\|{\mathbf{y}}^{*}({\mathbf{x}}^{k})-{\mathbf{y}}^{k+1}\|^{2}]+\frac{\zeta^{2}\ell\gamma}{2}\mathbb{E}[\|\nabla_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1};\xi)\|^{2}]+\frac{\alpha^{2}\ell d^{3}}{4\gamma}$
$\leq$ $\mathbb{E}[g({\mathbf{x}}^{k})-g({\mathbf{x}}^{k+1})]+\frac{\ell}{\gamma}\mathbb{E}[\|{\mathbf{y}}^{*}({\mathbf{x}}^{k})-{\mathbf{y}}^{k+1}\|^{2}]+\frac{\zeta^{2}\ell\gamma}{2}\mathbb{E}[\|\nabla_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]+\frac{\zeta^{2}\ell\gamma\sigma^{2}}{2B_{2}}+\frac{\alpha^{2}\ell d^{3}}{4\gamma},$
where the last inequality is from Assumption 5.1.
On the other hand, from Lemma A.3, by letting $\zeta=\frac{1}{2\ell\gamma}$ we have
$\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]$
$\leq$ $2\mathbb{E}[\|\nabla_{x}f_{\alpha}({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]+\frac{\alpha^{2}\ell^{2}(d+d^{\prime})^{3}}{2}$
$\leq$ $\frac{16}{3}\ell\gamma\mathbb{E}[g({\mathbf{x}}^{k})-g({\mathbf{x}}^{k+1})]+\frac{16}{3}\ell^{2}\mathbb{E}[\|{\mathbf{y}}^{*}({\mathbf{x}}^{k})-{\mathbf{y}}^{k+1}\|^{2}]$
$+\frac{2}{3}\mathbb{E}[\|\nabla_{\mathbf{x}}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]+\frac{2\sigma^{2}}{3B_{2}}+\frac{11}{6}\alpha^{2}\ell^{2}(d+d^{\prime})^{3}$
$\Rightarrow$ $\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]\leq 16\ell\gamma\mathbb{E}[g({\mathbf{x}}^{k})-g({\mathbf{x}}^{k+1})]+16\ell^{2}\mathbb{E}[\|{\mathbf{y}}^{*}({\mathbf{x}}^{k})-{\mathbf{y}}^{k+1}\|^{2}]$
$+\frac{2\sigma^{2}}{B_{2}}+\frac{11}{2}\alpha^{2}\ell^{2}(d+d^{\prime})^{3}.$
Taking summation of $k$ from $1$ to $K$ we have
$\frac{1}{K}\sum_{k=1}^{K+1}\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]$
$\leq$ $\frac{16\ell\gamma}{K}\mathbb{E}[g({\mathbf{x}}^{1})-g({\mathbf{x}}^{K+1})]+\frac{16\ell^{2}}{K}\sum_{k=1}^{K}\mathbb{E}[\|{\mathbf{y}}^{*}({\mathbf{x}}^{k})-{\mathbf{y}}^{k+1}\|^{2}]+\frac{2\sigma^{2}}{B_{2}}+\frac{11}{2}\alpha^{2}\ell^{2}(d+d^{\prime})^{3}$
$\leq$ $\frac{16\ell\gamma}{K}\mathbb{E}[g({\mathbf{x}}^{1})-\min_{\mathbf{x}}g({\mathbf{x}})]+\frac{16\ell^{2}}{K}\sum_{k=1}^{K}\mathbb{E}[\|{\mathbf{y}}^{*}({\mathbf{x}}^{k})-{\mathbf{y}}^{k+1}\|^{2}]+\frac{2\sigma^{2}}{B_{2}}+\frac{11}{2}\alpha^{2}\ell^{2}(d+d^{\prime})^{3}.$
Thus, by setting parameters as in Theorem 5.6 we have $\min_{k}\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]\leq\epsilon^{2}.$
On the other hand, since
$\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]=\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})-\nabla_{\mathbf{y}}f({\mathbf{x}}^{k},{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}]\leq\ell^{2}\mathbb{E}[\|{\mathbf{y}}^{k+1}-{\mathbf{y}}^{*}({\mathbf{x}}^{k})\|^{2}],$
similar to Lemma A.5 we have $\mathbb{E}[\|\nabla_{x}f({\mathbf{x}}^{k},{\mathbf{y}}^{k+1})\|^{2}]\leq\epsilon^{2}$ by setting $T=O(\kappa\log(\frac{\kappa}{\epsilon}))$ and $B_{1}=O(\epsilon^{-2}).$
Appendix B Experimental Setup
To recall the proposed Algorithm 2, we present a pipeline of the proposed Algorithm 2 in figure 1.
B.1 Single-Task experiments
Following MeZO (Malladi et al., 2023), we evaluate our approach on a range of classification and multiple-choice tasks. In this setting, training and testing are conducted on the same task.
B.1.1 Tasks
We use the following tasks for evaluating the fine-tuning capabilities of Bilevel-ZOFO in a single-task setting.
BoolQ (Clark et al., 2019):
A yes/no question-answering task where each question is paired with a paragraph that contains the answer.
CB (Wang et al., 2019):
The CommitmentBank task involves determining whether a given sentence in context entails, contradicts, or is neutral to a premise.
COPA (Roemmele et al., 2011):
The Choice of Plausible Alternatives (COPA) task requires selecting the most plausible cause or effect from two alternatives for a given premise.
ReCoRD: (Zhang et al., 2018)
The Reading Comprehension with Commonsense Reasoning Dataset (ReCoRD) is a cloze-style task where models must predict masked-out entities in text based on the surrounding context.
RTE (Wang, 2018):
The Recognizing Textual Entailment (RTE) task involves determining whether a given hypothesis is entailed by a provided premise.
SST2 (Wang, 2018):
The Stanford Sentiment Treebank (SST-2) task focuses on binary sentiment classification of sentences as positive or negative.
WiC (Pilehvar & Camacho-Collados, 2018):
The Word-in-Context (WiC) task involves determining whether the same word is used in the same sense in two different sentences.
WinoGrande (Sakaguchi et al., 2021):
A commonsense reasoning task where the goal is to resolve pronoun references in ambiguous sentences by identifying the correct antecedent.
WSC (Levesque et al., 2012):
The Winograd Schema Challenge (WSC) tests a model’s ability to resolve pronoun references in sentences, requiring commonsense reasoning.
SQuAD (Rajpurkar, 2016):
The Stanford Question Answering Dataset (SQuAD) is a reading comprehension task where models must answer questions based on a given passage of text.
B.1.2 PEFT Variants
We utilize three PEFT techniques—prompt-tuning (Lester et al., 2021), prefix-tuning (Li & Liang, 2021), and LoRA (Hu et al., 2022)—for lower-level training to evaluate bilevel-ZOFO across various conditions and resource constraints.
  1. 1.
    LoRA: For all single-task LoRA experiments, we set $r=8$ and $\alpha=16$.
  2. 2.
    Prefix Tuning: We use 5 prefix tokens across all experiments.
  3. 3.
    Prompt Tuning: We configure 10 soft prompt tokens for every experiment.
B.1.3 Hyperparameter Search
Given resource limitations, we focus on sweeping only the learning rate as the key hyperparameter. For MeZO and first-order PEFT experiments, we explore learning rates from the set $\{1e-2,1e-3,1e-4,1e-5,1e-6\}$. For Bilevel-ZOFO, we sweep both the upper-level and lower-level learning rates: $\text{lr}_{\text{upper}}\in\{1e-4,1e-5,1e-6\}$ and $\text{lr}_{\text{lower}}\in\{1e-2,1e-3,1e-4,1e-5\}$. We perform all experiments in tables 5 and 6 using three random seeds and report the average and standard deviation. We also set $\epsilon=1e-3$, following MeZO (Malladi et al., 2023).
B.1.4 Training
All experiments used a batch size of 8 and were conducted in bfloat16 precision on a single A6000 Ada 48GB GPU. MeZO was run for 10,000 steps, while FO and Bilevel-ZOFO methods were run for 5,000 steps. Our implementation builds upon MeZO’s codebase, and memory profiling as well as latency calculations are based on their framework.
For each task, 1000 examples are randomly sampled for training, 500 for validation, and 1000 for testing. For bilevel-ZOFO, the training set is split into upper-level and lower-level subsets with a 1:2 ratio. During each lower-level update, only the PEFT parameters are optimized, while in the upper-level step, the entire model is fine-tuned using zeroth-order gradient approximation. We set $\lambda=10000$ and perform 20 lower-level updates between each upper-level update for all bilevel-ZOFO experiments.
All experiments use the Adam optimizer (Kingma & Ba, 2015),including baselines and both lower-level and upper-level optimizers. No weight decay was applied, and the models were trained with a constant learning rate schedule. Batch size is set to $16$ for all experiments. We load all models in bfloat16. We find the best performing model based on validation loss and report test results from that checkpoint. We report the test accuracy or F1-score based on the test dataset being imbalanced or not.
We fix the memory budget of each step across bilevel-ZOFO and the baselines. We train zeroth-order methods for 10,000 steps, and bilevel-ZOFO and first-order methods for 5000 steps. We use A6000ada 48GPUs in our experiments. We load all models in bfloat16.
Figure 3 compares the number of parameters tuned by first-order PEFT methods that are used in our experiments.
Figure 3: Number of additional parameters PEFT methods introduced to each model.
B.2 Results
Figure 4: Training loss for the lower-level objective of the bilevel framework with Lora as the PEFT model.
Figure 4 presents the training loss for the lower-level objective of the bilevel framework with Lora as the PEFT model. As shown, consistent with the guarantees provided by our theoretical analysis, Bilevel-ZOFO converges.
Table 5 presents the test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B (Zhang et al., 2022) on a downstream task.
Trainer Mode BoolQ CB Copa ReCoRD RTE SST2 WIC WinoGrande WSC Average
MeZO ft $0.6927\pm 0.0660$ $0.7767\pm 0.1162$ $0.7000\pm 0.0289$ $0.6980\pm 0.0053$ $0.6587\pm 0.0271$ $0.8214\pm 0.0042$ $0.5543\pm 0.0146$ $0.5480\pm 0.0108$ $0.5054\pm 0.0056$ $0.6617\pm 0.0321$
lora $0.6860\pm 0.0012$ $0.7607\pm 0.0515$ $0.7200\pm 0.0058$ $0.7083\pm 0.0049$ $0.6755\pm 0.0110$ $0.8501\pm 0.0067$ $0.5549\pm 0.0057$ $0.5607\pm 0.0050$ $0.5570\pm 0.0000$ $0.6748\pm 0.0102$
prefix $0.6573\pm 0.0379$ $0.7945\pm 0.0309$ $0.7033\pm 0.0208$ $0.7047\pm 0.0010$ $0.6972\pm 0.0055$ $0.8218\pm 0.0127$ $0.5622\pm 0.0127$ $0.5370\pm 0.0137$ $0.5105\pm 0.1313$ $0.6654\pm 0.0285$
prompt $0.6260\pm 0.0056$ $0.5821\pm 0.0179$ $0.7067\pm 0.0058$ $0.7070\pm 0.0053$ $0.5415\pm 0.0063$ $0.7463\pm 0.0218$ $0.5574\pm 0.0048$ $0.5556\pm 0.0038$ $0.4654\pm 0.0618$ $0.6098\pm 0.0159$
average $0.6655$ $0.7285$ $0.7075$ $0.7045$ $0.6432$ $0.8099$ $0.5572$ $0.5503$ $0.5096$ $0.6529\pm 0.0217$
FO lora $0.7403\pm 0.0055$ $0.8512\pm 0.0412$ $0.7500\pm 0.0058$ $0.7206\pm 0.0035$ $0.7292\pm 0.0165$ $0.9258\pm 0.0032$ $0.6463\pm 0.0276$ $0.5806\pm 0.0055$ $0.6474\pm 0.0200$ $0.7324\pm 0.0143$
prefix $0.7300\pm 0.0035$ $0.8571\pm 0.0644$ $0.7167\pm 0.0115$ $0.7093\pm 0.0032$ $0.7136\pm 0.0110$ $0.8133\pm 0.0050$ $0.5387\pm 0.0050$ $0.5980\pm 0.0029$ $0.5705\pm 0.0294$ $0.6941\pm 0.0141$
prompt $0.7150\pm 0.0156$ $0.7142\pm 0.0714$ $0.7466\pm 0.0115$ $0.7163\pm 0.0063$ $0.6936\pm 0.0185$ $0.8016\pm 0.0779$ $0.5386\pm 0.0197$ $0.5980\pm 0.0090$ $0.5062\pm 0.0434$ $0.6700\pm 0.0306$
average $0.7284$ $0.8075$ $0.7378$ $0.7154$ $0.7121$ $0.8470$ $0.5745$ $0.5922$ $0.5747$ $0.6982\pm 0.0197$
Ours lora $0.7433\pm 0.0191$ $0.9167\pm 0.0103$ $0.7400\pm 0.0200$ $0.7183\pm 0.0031$ $0.7401\pm 0.0108$ $0.9331\pm 0.0020$ $0.6447\pm 0.0218$ $0.5903\pm 0.0058$ $0.6428\pm 0.0855$ $0.7410\pm 0.0209$
prefix $0.7340\pm 0.0095$ $0.8690\pm 0.0206$ $0.7267\pm 0.0153$ $0.7140\pm 0.0044$ $0.7304\pm 0.0091$ $0.8550\pm 0.0178$ $0.6317\pm 0.0282$ $0.5710\pm 0.0130$ $0.5810\pm 0.0338$ $0.7125\pm 0.0179$
prompt $0.7367\pm 0.0850$ $0.7679\pm 0.0644$ $0.7633\pm 0.0058$ $0.7257\pm 0.0153$ $0.6867\pm 0.0208$ $0.8335\pm 0.0779$ $0.6267\pm 0.0462$ $0.5900\pm 0.0173$ $0.5133\pm 0.1493$ $0.6938\pm 0.0536$
average $0.7380$ $0.8512$ $0.7433$ $0.7193$ $0.7191$ $0.8739$ $0.6344$ $0.5838$ $0.5790$ $0.7158\pm 0.0308$
Table 5: Single-Task Experiments on OPT-1.3B with 1000 samples. Values correspond to mean across three random seeds. FO: First-Order. FT: full-model fine-tuning.
Table 6 demonstrates the results for fine-tuning Llama2-7b (Touvron et al., 2023) on various classification and open-ended generation tasks.
Trainer Mode BoolQ ReCoRD SQuAD SST2 Average
MeZO ft 0.7915 ± 0.0516 0.7890 ± 0.0001 0.7737 ± 0.1634 0.8646 ± 0.0216 0.8047
lora 0.8020 ± 0.0014 0.7970 ± 0.0001 0.7412 ± 0.0013 0.8529 ± 0.0117 0.7983
prefix 0.7830 ± 0.0131 0.7905 ± 0.0007 0.7093 ± 0.0207 0.8364 ± 0.0010 0.7798
prompt 0.7787 ± 0.0049 0.7935 ± 0.0007 0.7014 ± 0.0451 0.8246 ± 0.0216 0.7746
FO lora 0.8420 ± 0.0104 0.7920 ± 0.0053 0.8197 ± 0.0043 0.9557 ± 0.0007 0.8524
prefix 0.7783 ± 0.0021 0.8013 ± 0.0012 0.7946 ± 0.0419 0.9243 ± 0.0053 0.8246
prompt 0.8083 ± 0.0142 0.8023 ± 0.0074 0.7805 ± 0.0633 0.9284 ± 0.0072 0.8299
[…]
Table 6: Single-Task Experiments on Llama2-7B with 1000 samples. Values correspond to mean and std across three random seeds. FO: First-Order. FT: full-model fine-tuning
B.3 Memory Profiling and Wall Clock Time Analysis
Figure 5 demonstrates the memory profiling of Bilevel-ZOFO, MeZO and First-order prefix tuning on four different tasks. Memory consumption of MeZO and first-order PEFT methods varies across tasks, with one occasionally surpassing the other. Each lower-level update in our method matches that of the corresponding PEFT method. Similarly, each upper-level update requires the greater memory usage between MeZO and PEFT under comparable settings. As a result, the total memory requirement of our method corresponds to the maximum memory usage of the PEFT and MeZO experiments. Nonetheless, as demonstrated in Table 5, our method outperforms both PEFT and MeZO on average.
We also present a wall-clock time analysis of bilevel-ZOFO compared to the baseline. As shown in Table 7, similar to MeZO (Malladi et al., 2023), we observe that zeroth-order steps exhibit higher latency compared to first-order steps. The results indicate that our bilevel-ZOFO achieves comparable delays to the FO-PEFT method while significantly reducing step duration compared to MeZO. Moreover, as highlighted in Table 2, bilevel-ZOFO outperforms both methods on average.
Figure 5: Memory consumption of MeZO and first-order PEFT methods varies across tasks, with one occasionally surpassing the other. Our Bilevel-ZOFO method demonstrates comparable memory usage to both baselines. Values correspond to memory usage for fine-tuning OPT1.3b (Zhang et al., 2022) on each task using a batch size of 8 and on a singel A6000ada 48GB GPU.
Table 7: Wallclock time per step of different training methods when finetuning OPT1.3b. The values are measured on a single A6000ada 48GB GPU. The wallclock time is averaged over 3 different runs that produced the values of Table 2. We use a batch size of 8 for all experiments.
Task MeZO FO Prefix-Tuning Bilevel-ZOFO (Prefix)
Copa 0.299 0.127 0.135
MultiRC 0.622 0.474 0.502
WSC 0.278 0.120 0.164
B.4 Sensitivity Analysis
In this section, we provide experimental results that prove Bilevel-ZOFO mitigates the sensitivity of MeZO to hard prompts. First, similar to Table 5 of the MeZO (Malladi et al., 2023) paper, we experiment with different choices of hard prompts for both MeZO and our Bilevel-ZOFO. Table 8 shows the results for tuning Opt 1.3b on SST2 and COPA. We can see that our method effectively mitigates the sensitivity of MeZO to hard prompts. The difference between the results with and without a simple hard prompt in our experiment is much less than MeZO’s
Method Experiment With Prompt (%) Without Prompt (%) Difference
MeZO SST-2 89.6 51.9 -38.6
COPA 70.0 54.8 -15.2
Bilevel-ZOFO SST-2 93.3 92.9 -0.4
COPA 76.66 73.6 -3.06
Table 8: Prompt Sensitivity Comparison for MeZO and Bilevel-ZOFO. Bilevel-ZOFO effectively mitigates the sensitivity of MeZO to hard prompts.
To also validate that the improved results are not because of tuning more parameters, we conducted an experiment on COPA using OPT1.3B and compared Bilevel-ZOFO to a two-stage pipeline that tunes the same number of parameters. First, we performed first-order prompt tuning for a fixed number of steps (same as the number of lower-level updates in bilevel-ZOFO), followed by additional tuning using ZO for the same number of iterations as the upper level updates in bilevel-ZOFO (A two-stage pipeline). As shown in Table 9, even with extensive hyperparameter tuning, the second stage does not improve the results achieved after the first stage and is highly likely to decrease performance. Our method, however, improves performance when using the same number of steps in the upper and lower levels, respectively. The bilevel structure makes the trained prompts dynamically optimal for the full ZO fine-tuning and reaches an accuracy of 76.66.
The observed performance drop after the second stage is indeed counter-intuitive at first glance. However, it is a limitation of MeZO as it approximates gradients. While further fine-tuning intuitively should improve performance, the inherent noise in gradient approximation can lead to suboptimal updates. This observation is consistent with the fact that MeZO typically requires a significant number of iterations to converge. This is a key contribution of our work: Our approach addresses MeZO’s challenges, such as sensitivity to hard prompts and long convergence times, while outperforming both MeZO and PEFT and maintaining similar memory efficiency. The intuition behind why our method is effective in enhancing both MeZO’s full-model tuning and PEFT is in the nested bilevel structure. This structure encodes more information (as reflected in the training method) from the prompt tuning stage than only treating it as a first stage, thereby providing better guidance for MeZO. In contrast, our bilevel method effectively addresses the issues of MeZO and demonstrates improved performance over both MeZO and the PEFT baseline, even with the same number of ZO iterations. See  (Shirkavand et al., 2024) for a more detailed analysis of why a bilevel-method is better than a two-staged pipeline.
Method Experiment (COPA) Accuracy (%)
Two-Stage Pipeline After Stage 1 74.33
After Stage 2 (lr $0.001$) 51.66
After Stage 2 (lr $0.0001$) 70.33
After Stage 2 (lr $0.00001$) 72.66
After Stage 2 (lr $0.000001$) 74.33
Bilevel-ZOFO - 76.66
Table 9: Comparison of Bilevel-ZOFO with a two-staged pipeline.
The training loss curves for both stages of a two-stage approach and our bilevel framework are provided in Figure 6. When running MeZO in the second stage, the training loss exhibits oscillations and does not show improvement within 500–1000 iterations. This behavior is consistent with findings in the original MeZO (Malladi et al., 2023) paper, which notes that MeZO typically requires much longer to converge—on the order of 100k iterations. The oscillatory behavior observed within the shorter training duration is not surprising due to gradient approximation errors.
(a) First-Stage PEFT
(b) Second-Stage MeZO
(c) Bilevel
Figure 6: The training loss curves for both stages of a two-stage approach (a and b) and our bilevel framework (c).
B.5 Multi-task experiments
In this section we explain the experimental details of mutil-task experiments.
B.5.1 Meta-Tasks
Following the methodology of Min et al. (2022), we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings outlined in MetaICL (Min et al., 2022): classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT (Ye et al., 2021) and UNIFIEDQA (Khashabi et al., 2020), comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification, question answering, and natural language understanding, all in English. Table 10 shows the number of tasks in each training and testing meta-learning setting and the total number of examples in each training task.
Meta-train Setting # tasks # examples Target Setting # tasks
Classification 43 384,022 Classification 20
Non-Classification 37 368,768
QA 37 486,143 QA 22
Non-QA 33 521,342
Table 10: Details of four different meta-learning settings. Each row indicates meta-training/target tasks for each setting. There is no overlap between the training and test tasks.
See Tables 14 and 15 of MetaICL (Min et al., 2022) for a list of all sub-tasks.
B.5.2 Baselines
We use GPT2-Large (Radford et al., 2019) as the base model for these experiments.We compare our method against several baseline approaches:
  • MetaICL (Min et al., 2022): A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, ${(a_{1},b_{1}),\dots,(a_{k},b_{k})}$, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
  • Zero-shot: This method uses the pretrained language model (LM) without any tuning, performing zero-shot inference without any demonstration examples.
  • In-context Learning (ICL): This method uses the pretrained LM with in-context learning by conditioning on a concatenation of $k$ demonstration examples and 1 actual test sample similar to MetaICL.
We sample 768 examples from each training sub-task. We use these samples to train MetaICL in their original setting for 30,000 steps. This includes learning rate of $51e-$ 5, batch size of $1$ on $8$ GPUs, 8-bit Adam optimizer and fp16 half precision. See MetaICL(Min et al., 2022) for full details. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we do not include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  (Malladi et al., 2023), we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements. We use Adam optimizer and a learning rate of $61e-$ 6 for both upper and lower-level training. We employ a batch size of $4$ and train on a single rtx6000ada GPU.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on $4$ demonstration samples from each class of each test sub-task. Similar to (Min et al., 2022), we report Macro-averaged F1 as the evaluation metric.